### v6

@@ Abstract
Automatic breast cancer classification technology in histopathology images plays a crucial role in accurate diagnosis and treatment planning. Recently, histopathology image classification approaches based on ResNet architecture have been widely proposed as they vastly improve the classification accuracy using skip connection to prevent vanishing gradient issues by combining low-level and high-level feature information. However, traditional ResNet architecture suffers from challenges such as data imbalance and interpretability, which still require cross-domain knowledge and collaboration among medical experts. This study proposes a new breast cancer classification method using a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) approach to address data imbalance and interpretability challenges effectively. The proposed approach combines a pre-trained deep ResNet50 method and a lightweight attention mechanism to accomplish the classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers based on LeakyReLU and ReLU activation functions to enhance the ability of feature learning. The proposed DALAResNet50 approach has been evaluated on breast cancer histopathology images derived from BreakHis Datase at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. The DALAResNet50 approach was comprehensively assessed with popular deep learning approaches such as SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models. The reported results of DALAResNet50 outperform the compared approaches regarding precision, accuracy, recall, F1 score, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and unbalanced breast cancer datasets.
@@ I Introduction
Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2]. Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer. However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer. In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer. Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis. Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step. Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency. These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3]. This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7]. One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data, such as medical images, with a grid-like structure. The concept of CNN was first introduced back in 1998 [8]. The concept describes a method for efficiently classifying image data. CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to remove necessary features from different categories of pictures so that the other categories can be classified accurately based on those features. CNNs have three main layers: convolutional, pooling, and fully connected. These layers work together to make CNNs a powerful tool for classifying medical images. The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig. 1. However, despite the considerable achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data. To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures. This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images. The critical contributions of the work can be summarized as follows:
@@
(a) Proposing a lightweight attention mechanism network that adaptively adjusts channel feature extraction to improve feature representation.
@@
(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques. Specifically, in the fully connected network component of the attention mechanism module, LeakyReLU [10] and ReLU [11] activation functions were utilized. LeakyReLU was selected for its ability to mitigate the gradient vanishing problem. In contrast, ReLU was employed to enhance the representation of image features. This choice was instrumental in better capturing multi-scale spatial patterns.
(c) In this paper, test images with different magnifications and verify that the network is invariant to image scales and performs better than other state-of-the-art methods.
@@
(d) The model was thoroughly evaluated by analyzing loss. The results show stable performance in the model’s later training phases, minimizing overfitting and exhibiting strong generalization for unfamiliar data.
This paper is organized as follows:
Section 2 provides an overview of the current state of the art in applying deep learning to breast cancer diagnostic techniques.
Section 3 details the proposed breast cancer classification method based on dual activation lightweight attention ResNet50.
@@
Section 4 presents the dataset and implementation details of the experimental task and compares the classification results with other state-of-the-art networks.
@@
Section 5 comprehensively evaluates the model’s performance through loss and accuracy analysis.
Finally, Section 6 provides a summary and conclusions.
@@ II Related works
Currently, two main methods are primarily used for cancer tissue histopathology image classification: traditional manual feature extraction and deep learning-based methods. Traditional methods rely on manually extracting image features and then using classifiers such as SVM or Random Forest for recognition; however, this method has shortcomings, such as a time-consuming feature extraction process, the requirement of domain knowledge, and the challenge of securing high-quality features [12]. In contrast, methods based on deep learning extract features through models like convolutional neural networks, which is more efficient [13]. Deep learning-based cancer tissue histopathology image classification primarily encompasses three methods: CNN-based breast cancer classification, transfer learning-based breast cancer classification, and attention mechanism-based breast cancer classification. The following section delves into these techniques in detail.
@@ II-A CNN
CNN [8], one of the most prevalent deep learning techniques, is apt for tasks involving image processing [14]. Through convolutional, pooling, and fully connected layers, CNNs are highly effective in learning image features and classifying them [15]. In breast cancer histopathology image classification, CNN models with various structures, such as AlexNet, Inception, VGGNet, DenseNet, ResNet, etc., can be trained on large datasets to learn features and classify breast cancer.
AlexNet, a deep convolutional neural network introduced by Alex et al. in 2012 [16], marked a pivotal breakthrough in the ImageNet image classification competition, setting a landmark in deep learning and CNNs. Omonigho et al. [17] employed a modified AlexNet with augmentation techniques on the MIAS mammography database for breast cancer detection, achieving an impressive 95.70% accuracy. This surpassed traditional methods, though parameter optimization and validation on unbalanced datasets require further exploration.
The Inception model is a neural network architecture introduced by Google that aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks. The hybrid model based on CNN-Inception-V4 was used by Nazir et al. [18]. It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
VGGNet, a deep neural network architecture developed by researchers at Oxford University, performs well in image classification tasks. However, it is unsuitable for embedded devices or resource-constrained environments because of its high computational and memory requirements due to its considerable depth and number of parameters. Nedjar et al .[19] proposed a visualization method for explaining and understanding the decision-making process of VGG19 in classifying breast biopsy images, which helps to increase the trust and acceptance of medical professionals while providing a method for evaluating and Enhancing the efficiency of CNNs.
DenseNet is a deep neural network architecture developed by Gao Huang et al. [20]. It is known for its unique connectivity pattern. Each layer is densely connected to all previous layers, thus enabling the effectiveness of information flow and gradient propagation in deep networks. Jiménez Gaona et al. [21] introduced a novel approach to breast cancer detection and classification using the DenseNet architecture, which has advantages over other pre-trained CNN methods in solving problems such as gradient vanishing and reducing the number of parameters.
ResNet, a deep convolutional neural network architecture introduced by He et al. in 2015 [9], is particularly notable for its solution to the vanishing gradient issue in deep neural networks. This enables the development of intense neural networks, leading to substantial improvements in tasks like image categorization. Khikani et al. [22] proposed an enhanced capsule network using Res2Net blocks and additional convolutional layers to extract multi-scale features, and the results showed that their model outperformed previous deep-learning methods for breast cancer classification. However, ResNet-50 is prone to problems in handling unbalanced datasets. Given the significant superiority of the ResNet50 model over other methods in terms of accuracy, this paper has decided to modify the ResNet50 model to address these specific challenges.
@@ II-C Attention mechanism
In medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique. It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions. By concentrating on crucial features within these images, the attention mechanism dramatically enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes. In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images crucial for accurately distinguishing between malignant and benign tissues. This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For instance, Vardhan et al. [35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer. The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis. This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions. Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms. This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism. The insights gleaned indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and boost the interpretability of the model. However, this also correspondingly increases the following difficulties:
@@ II-C1 Limitations in global feature extraction
Although the attention mechanism is designed to enhance model performance by focusing on critical parts of the image, thereby playing a positive role in strengthening the model’s ability to process global information, it may still face challenges in some implementations or specific application scenarios, unable to fully capture the international contextual details of the entire image.
@@ II-C2 Insufficiency in feature priority precision
Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information. This approach risks overlooking vital features essential for precise classification, undermining the model’s ability to highlight and utilize discriminative patterns within the data effectively.
@@ II-C4 Lack of feature adaptability
This study designed an innovative lightweight attention module to overcome the identified difficulties. This design incorporates adaptive average pooling, a two-layer fully connected network for channel recalibration, dual activation functions of LeakyReLU [10] and ReLU [11], and element-wise multiplication. Adaptive average pooling tackles the limitations in global feature extraction, whereas the two-layer fully connected network for channel recalibration optimizes feature priority precision. The combination of LeakyReLU and ReLU effectively addresses the inefficiencies of a single activation function, enhancing activation function efficiency, and element-wise multiplication improves feature adaptability. These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models.
@@ III-A Transfer ResNet50
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table I. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before. This design breakthrough allowed for achieving exceptional performance. However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building. Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
@@
Note: The term ”1000-d FC” in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
@@
Feature
DALAResNet50 model
Standard ResNet model
Adaptive Average Pooling
Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
Not available
Channel Recalibration
A two-layer fully connected network for channel recalibration optimizes ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
Not available
Activation Functions
The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, non-zero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers. This combination maintains the model’s learning capability, potentially enhancing the training process.
ReLU only
Element-wise multiplication
Element-wise multiplication, as denoted by ’Scale’ in Fig. 3, enhances ”Feature Adaptability” by allowing dynamic, non-linear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
Not available
Fig. 2: Flowchart for standard ResNet model
Fig. 3: Flowchart for DALAResNet50 model

### v8

@@ Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets.
@@ I Introduction
Breast cancer has become a leading cause of mortality, particularly among women, with its incidence rising continuously, raising global health concerns [1, 2]. Early, accurate diagnosis and treatment are pivotal in decreasing the mortality rate associated with breast cancer. Yet, medical practices face challenges such as insufficient healthcare resources and human error in interpretation, which impede the timely and precise detection and diagnosis of breast cancer. Against this backdrop, intelligent diagnostic assistance technologies have emerged as innovative solutions in the battle against breast cancer. Recent advances in deep learning have revolutionized breast cancer diagnosis, surpassing the limitations of traditional classification and recognition algorithms by integrating feature extraction, selection, and classification into a unified step. Deep learning automates feature extraction and classification through neural networks, reducing redundancy and enhancing efficiency, thus significantly improving diagnostic accuracy and workflow [3, 4, 5, 6, 7]. A pivotal innovation in this domain is developing and applying Convolutional Neural Networks (CNNs), specialized deep learning models adept at processing data, such as medical images, with a grid-like topology. In 1998, LeCun et al. [8] introduced Convolutional Neural Networks (CNNs), which progressively extract abstract features from images through multiple levels of convolutional and pooling operations. This process enables CNNs to achieve classification based on these extracted features. CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images. The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes, as illustrated in Fig. 1. Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data. This paper proposed a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model designed to overcome the challenge of imbalanced data in CNN architectures. This innovative approach efficiently manages uneven category distributions and enhances the detection of malignant breast cancer in medical images. The primary contributions of this work are summarized as follows:
@@
(a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation. This approach optimizes the network’s ability to discern relevant features for breast cancer classification.
@@
(b) Integrating the attention mechanism with ResNet50 [9] creates a feature-rich model, achieving more efficient feature extraction by combining the strengths of both techniques. Specifically, within the attention mechanism’s fully connected network component, LeakyReLU [10] and ReLU [11] activation functions are employed. LeakyReLU is chosen for its capability to alleviate the issue of vanishing gradients; ReLU is used to improve the representation of image features, facilitating the capture of multi-scale spatial patterns more effectively.
(c) The paper tests images at various magnifications, demonstrating the network’s scale invariance and superior performance over other contemporary methods.
@@
(d) The model undergoes comprehensive evaluation, focusing on analyzing loss. The findings indicate stable performance during the later stages of training, effectively minimizing overfitting and demonstrating strong generalization capabilities for new data.
Organization of the paper:
Section II reviews the latest advancements in applying deep learning for breast cancer diagnosis.
Section III proposed the DALAResNet50 model approach for classifying breast cancer histopathology images.
@@
Section IV outlines the dataset and implementation specifics of the experimental setup and compares the classification outcomes with those of other leading networks.
@@
Section V conducts an in-depth performance evaluation of the model through loss and accuracy analysis.
Section VI provides a summary and conclusions.
@@ II Related works
Cancer tissue histopathology image classification predominantly utilizes traditional manual feature extraction and deep learning-based methods. Traditional approaches depend on manually extracting features from images and employing classifiers such as SVM (Support Vector Machine) or Random Forest for recognition. However, these methods encounter limitations, such as the need for intensive feature extraction, a requirement for substantial domain knowledge, and challenges in acquiring high-quality features [12]. In contrast, deep learning-based methods, mainly through models like Convolutional Neural Networks (CNNs), offer a more efficient feature extraction process [13]. Specifically, deep learning approaches to cancer tissue histopathology image classification include CNN-based, transfer learning-based, and attention mechanism-based breast cancer classification. Each of these methods leverages the inherent strengths of deep learning to improve classification performance, which will be explored in detail in the subsequent section.
@@ II-A CNN
CNN, one of the most widespread deep learning techniques, excels in image processing tasks [14]. They utilize convolutional, pooling, and fully connected layers to learn and classify image features effectively [15]. In breast cancer histopathology image classification, CNN models like AlexNet, Inception, VGGNet, DenseNet, and ResNet have demonstrated substantial capabilities when trained on large datasets.
AlexNet, introduced in 2012 by Alex et al. [16], represented a significant milestone in ImageNet competition history, showcasing the potent capabilities of CNNs. Modified versions of AlexNet, such as the one by Omonigho et al. [17] on the MIAS mammography database, achieved notable accuracy, indicating superior performance over traditional methods. However, challenges remain in parameter optimization and application to imbalanced datasets.
The Inception model, developed by Google, addresses computational and parameter efficiency in neural networks, making it ideal for image classification tasks. Nazir et al.’s hybrid CNN-Inception-V4 model [18] demonstrated high accuracy and sensitivity, though it faced limitations in specificity and computational efficiency considerations.
VGGNet, crafted by Oxford University researchers, excels in image classification but is less suited for resource-constrained environments due to its intensive computational and memory demands. Nedjar et al. [19] presented a visualization method for VGG19’s decision-making process in breast biopsy image classification, enhancing medical professionals’ trust and CNN efficiency.
DenseNet [20], known for its efficient information and gradient flow through dense connectivity, offers advantages in overcoming challenges like gradient vanishing. Jiménez Gaona et al. [21] leveraged DenseNet for breast cancer detection, showing its superiority in reducing parameters and improving performance compared to other CNN architectures.
ResNet, introduced by He et al. in 2015 [9], addresses deep neural networks’ vanishing gradient problem, allowing for the construction of deeper networks. An enhanced capsule network using Res2Net blocks, proposed by Khikani et al. [22], demonstrated improved performance in breast cancer classification, although ResNet-50 faces challenges with imbalanced datasets. This paper presents modifications to the ResNet50 model to tackle these issues, given its accuracy advantage over other methods.
@@ II-C Attention mechanism
In medical imaging, the attention mechanism [32] in deep learning has emerged as a revolutionary technique. It allows models to allocate varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This selective focus enables the models to prioritize and scrutinize the most relevant areas for diagnosing diseases, detecting anomalies, or monitoring the progression of medical conditions. By zeroing in on essential features within these images, the attention mechanism significantly boosts the accuracy and efficiency of automated medical image analysis, contributing to early detection and enhanced patient outcomes. Within breast cancer image classification, incorporating the attention mechanism encourages the model to discern critical and distinctive features in breast cancer images that are pivotal for accurately differentiating between malignant and benign tissues. This strategy can elevate the model’s performance, diminish reliance on global image feature extraction, and enhance the precision in pinpointing vital breast cancer regions [34].
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For example, Vardhan et al. [35] developed a breast cancer segmentation system utilizing Convolutional Neural Networks (CNNs) with interpretable artificial intelligence and the attention mechanism to enhance breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation employing a modified DenseNet architecture with an attention mechanism to refine the CNN’s performance on the BreakHis dataset for high-precision breast cancer classification. This method was evaluated on the BreakHis dataset using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep transfer learning and visual attention mechanisms to increase the accuracy and efficiency of real-world clinical diagnoses. Based on deep transfer learning with attention mechanisms, this approach utilizes transfer learning to tackle the overfitting problem and integrates an attention module to concentrate on characterizing breast lesions. Wang et al.’s work achieves high-precision results that surpass other state-of-the-art breast image classification algorithms.
This section summarizes recent scholarly research and exploration of breast cancer classification, emphasizing the use of the attention mechanism. The insights indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and increase the model’s interpretability. However, this also correspondingly increases the following difficulties:
@@ II-C1 Limitations in global feature extraction
Although the attention mechanism is designed to improve model performance by focusing on critical sections of the image, thereby playing a positive role in enhancing the model’s ability to process global information of the image, it may still encounter challenges in some implementations or specific application scenarios, failing to grasp the entire image’s global contextual details fully.
@@ II-C2 Insufficiency in feature priority precision
Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information. This approach risks overlooking vital features essential for precise classification, compromising the model’s effectiveness in identifying and leveraging distinctive patterns in the data.
@@ II-C4 Lack of feature adaptability
This study designed an innovative lightweight attention module to overcome the identified difficulties. This design incorporates adaptive average pooling, a two-layer fully connected network for channel recalibration, dual activation functions of LeakyReLU [10] and ReLU [11], and element-wise multiplication. Adaptive average pooling tackles the limitations in global feature extraction, whereas the two-layer fully connected network for channel recalibration optimizes feature priority precision. The combination of LeakyReLU and ReLU effectively addresses the inefficiencies of a single activation function, enhancing activation function efficiency, and element-wise multiplication improves feature adaptability. These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models. Moreover, this paper presents a comparative analysis between the proposed DALAResNet50 model and the standard ResNet model, as illustrated in Table 2.
@@ III-A Transfer ResNet50
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table I. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks. This design breakthrough allowed for achieving promising performance. Therefore, the transfer learning strategy based on the pre-trained ResNet50 model is adopted in this paper, considering the efficient utilization of data and computing resources for model building. The benefit of transfer learning is the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities. As shown in Fig. 2, the proposed approach includes a histopathology breast cancer database and components for preprocessing and data augmentation, transfer learning based on ResNet50, and classification based on an improvised dual-activated fully connected layer. The proposed approach involves the following stages for breast cancer classification using histopathology images:
@@
Note: The term ”1000-d FC” in the pooling layer refers to a fully connected layer consisting of 1000 neurons.
@@
Feature DALAResNet50 model Standard ResNet model
Adaptive Average Pooling Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding. Not available
Channel Recalibration A two-layer fully connected network for channel recalibration optimize ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification. Not available
Activation Functions The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, nonzero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers. This combination maintains the model’s learning capability, potentially enhancing the training process. ReLU only
Element-wise multiplication Element-wise multiplication, as denoted by ’Scale’ in Fig. 3, enhance ”Feature Adaptability” by allowing dynamic, nonlinear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information. Not available
Fig. 2: The framework of DALAResNet50 model
Fig. 3: Comparison of conventional ResNet50 and DALAResNet50 model: (a) ResNet50 model (b) DALAResNet50 model